OpenStudy (anonymous):

FREE MEDAL FOR ANSWER please solves these for me and I will give medal....please please The radius of the base of a right circular cylinder is 3.72 ft. The height of the cylinder is 7.5 ft. Find the surface area of the cylinder.

OpenStudy (anonymous):

surface area of the cylinder is 262.3

OpenStudy (anonymous):

215.136 sq ft

OpenStudy (anonymous):

Surface area of a cylinder=\[2 h \pi r+2 \pi r^2=2 (7.5 )\pi (3.2)+2 \pi (3.2)^2=215.136 \]

OpenStudy (anonymous):

there are two bases on a cylinder

OpenStudy (anonymous):

The area of one base is pi r^2. The area of two bases is 2 pi r^2. That is what the second term of the area expression above is. The surface area of the two bases is 150.796 and the area of the cylinder missing the bases, pi d h or pi 2r h, is 64.3398.